A uniform magnetic field, B= 3.0 x 10-4T in the +x direction. A proton (q=-+e) shoots through the field in the +y direction with a speed of 5.0 x 106m/s.

(a) Find the magnitude and direction of the force on the proton.

(b) Repeat with the proton replaced by an electron.


1
Expert's answer
2021-04-28T07:43:21-0400

To be given in question

Magnetic force (B)="3\times10^{-4}T"

Charge of proton (q)="+1.6\times10^{-19}C"

Velocity (v)="5\times10^6m/sec"

"\theta=90°"

To be asked in question

(a) magnetic force (for proton)=?

(b) magnetic force (for electron)=?

We know that

(a) magnetic force "( \overrightarrow F)=q(\overrightarrow{v}\times \overrightarrow{B})"

"F=qvBsin\theta""\hat{n}"

sin90°=1

Put values

"F =qvB\hat{n}"

"\hat{n}" direction is "( \overrightarrow v\times \overrightarrow B)" direction



"F={1.6\times10^{-19}\times5\times10^6\times3\times10^4}\hat{ n}" "F=2.4\times10^{-8} (\hat{n})N"

"\hat{n}" direction is "-\hat{k}" direction

Magnitude of magnetic force

"F=2.4\times 10^{-8}""-(\hat{k})"

Magnetic force direction "-\hat{k}"

Negative (-) z- axis


(b) for electron magnetic force

"\overrightarrow F=q( \overrightarrow v\times \overrightarrow B)"

"F=qvBsin\theta(\hat{n})"

"\hat{n}" direction is "( \overrightarrow v\times \overrightarrow B)" direction

Put value


"F={-1.6\times10^{-19}\times5\times10^6\times3\times10^4}(\hat{n})" "F=-2.4\times10^{-8}(\hat n)N"

"\hat{n}" direction is "-\hat{k}" But charge is"q=-1.6\times10^{-19}C" then direction is"-(-\hat n)=\hat{n}"

Resultant "\hat{n}" direction is"+\hat{k}" direction

"F=2.4\times 10^{-8}(\hat n )N"

Magnitude of magnetic force

"F=2.4\times 10^{-8}N" "(+\hat{k})"

magnetic force direction

"(+\hat{k})"

Positive (+ )z-axis